There are 210 students doing the morning exercise. Kinder, Grade Five, and grade six. There are 15 more grade five children than Kinder and 45 less grade six children than kinder. How many children from each grade are exercising?
Answer by nerdybill(7384)

There are 210 students doing the morning exercise. Kinder, Grade Five, and grade six. There are 15 more grade five children than Kinder and 45 less grade six children than kinder. How many children from each grade are exercising?
.
Let k = number kinder children
then from "There are 15 more grade five children than Kinder "
k+15 = grade five
and from "45 less grade six children than kinder"
k-45 = grade six
.
since there are a total of 210:
210 = k + (k+15) + (k-45)
210 = k + k + 15 + k - 45
210 = 3k - 30
240 = 3k
80 = k (number of kids in kindergarten)
.
number of kids in grade five:
k+15 = 80+15 = 95
.
number of kids in grade six:
k-45 = 80-45 = 35